### v1

# Knowledge-Centric Templatic Views of Documents

Abstract
Authors seeking to communicate with broader audiences often compose their ideas about the same underlying knowledge in different documents and formats – for example, as slide decks, newsletters, reports, brochures, etc. Prior work in document generation has generally considered the creation of each separate format to be different a task, developing independent methods for generation and evaluation. This approach is suboptimal for the advancement of AI-supported content authoring from both research and application perspectives because it leads to fragmented learning processes, redundancy in models and methods, and disjointed evaluation. Thus, in our work, we consider each of these documents to be templatic views of the same underlying knowledge, and we aim to unify the generation and evaluation of these templatic views of documents. We begin by introducing an LLM-powered method to extract the most important information from an input document and represent this information in a structured format. We show that this unified representation can be used to generate multiple templatic views with no supervision and with very little guidance, improving over strong baselines. We additionally introduce a unified evaluation method that is template agnostic, and can be adapted to building document generators for heterogeneous downstream applications. Finally, we conduct a human evaluation, which shows that humans prefer 82% of the downstream documents generated with our method. Furthermore, the newly proposed evaluation metric correlates more highly with human judgement than prior metrics, while providing a unified evaluation method.
Figure 1: Visualization of our method. Given an input document, we prompt the LLM to generate a Structure Unified Representation (SURe). We can use the SURe to prompt the model to generate a templatic view of the input document. We then evaluate the generations using our unified evaluation.
1 Introduction
Sharing information in multiple formats is an important part of communication in a variety of domains, as it allows for knowledge to be disseminated to a wider audience. A product manager may need to create a requirements document, a product pitch deck, and a product announcement newsletter. Likewise, a person on the job market may create a resume, a cover letter, and a personal website. We consider these documents to be templatic views of the same underlying knowledge.
This is also true for the scientific domain, in which researchers need to create documents in multiple formats to effectively disseminate their work, – such as through conference talks, social media posts, poster presentations, and non-technical blog posts. Sharing knowledge in multiple formats broadens its audience and can help bridge the information gap between domain experts, researchers in adjacent fields, and even the general public – leading to greater understanding, new collaboration and accelerated progress in the field. With the significant increase in the pace of publication in recent years Bornmann and Mutz (2014), it has become essential for scientists to engage with the community and share their work in more compact, efficient, or easier to digest formats.
Past work on document generation has focused on developing methods of generation and evaluation for each specific document type individually Fu et al. (2021); Qiang et al. (2016); Chandrasekaran et al. (2020). However, tailoring methods to each template is time consuming and limiting, resulting in methods that are usable only for each intended template that cannot be compared across template types with specialized evaluation metrics. Our goal is to unify the tasks of generating and evaluating templatic views of documents. By doing so, we aim to allow system designers to adapt to new document types, templates and domains easily and efficiently, and to allow researchers to compare approaches more flexibly and fairly.
We begin by setting a few desiderata for our system for templatic view generation; it must be (1) unified, (2) adaptable, and (3) require minimal supervision. In order to design such a system to generate each of these templates, the method must be able to create structure and follow style norms associated with each template. We thus introduce an LLM-powered approach to generate a Structure Unified Representation (SURe) that represents the most important knowledge of an input document in a structured format. We demonstrate that this SURe can be used with no supervision and with very little prompted guidance about the output formats, to generate multiple templatic views of a document.
Meanwhile, automatic evaluation is crucial for researchers and developers to iterate more quickly than they otherwise would be capable of by relying solely on human evaluation. However, because in the past each templatic view has been tackled independently, previous work has used different evaluation metrics for each task, such as BERTScore for summarization Zhang* et al. (2020), Slide Level ROUGE for slide decks Qiang et al. (2016), or even a custom-trained, model-based metric for posters Wang et al. (2015). In this paper, we introduce Template Agnostic Evaluation (TAE) – a unified precision-recall style method of automatic evaluation that is template agnostic. Importantly, this evaluation framework is highly customizable, allowing users to easily integrate existing text-based metrics from the literature into its formulation and tailor its function to their specific use case.
We evaluate our unified approach to templatic view generation and evaluation on 3 types of documents: slides, posters, and blog posts. In each of these, the inputs are scientific articles and the outputs are expected to be documents of the corresponding types. We use the DOC2PPT Fu et al. (2021), Paper-Poster Qiang et al. (2016), and LongSumm Chandrasekaran et al. (2020) datasets, respectively. We show that our method of generating a Structured Unified Representation leads to improvements in performance across each tasks, demonstrating that LLMs are capable of generating templatic views with structure-augmented prompts, despite having no supervision. We additionally conduct a human evaluation to validate both our document generation method and evaluation metric. We show that humans prefer the output yielded by our method 82% of the time and that our evaluation metric correlates more highly with human preference than other popular metrics, such as ROUGE and BERTScore Lin (2004); Zhang* et al. (2020).
While the focus of this paper is the transformation of scientific documents, this choice was primarily driven by the availability of publicly available data. We note that nothing about our unified framework is domain specific, and we believe our approach for the creation and evaluation of templatic views of documents is easily adaptable to novel domains, and document types. More generally, we believe our work is a step towards the broader and more challenging goal of AI-assisted authoring support, and future research in reasoning over and generation of long documents.
2 Related Work
There are several areas of related research in NLP that are relevant to the problems of document transformation and evaluation.
Document summarization has been explored in a number of domains, including news See et al. (2017), literature Sciré et al. (2023), law Deroy et al. (2023), and dialogue Chen et al. (2021). In the scientific domain, summarization of scientific papers has taken the form of long form summaries Chandrasekaran et al. (2020), abstract generation Cohan and Goharian (2015), conference talks Lev et al. (2019), and query based summaries Fok et al. (2023). These summaries can be either extractive Sefid and Giles (2022) or abstractive Chandrasekaran et al. (2020).
Although the tasks of slide and poster generation have generally been considered separate from scientific summarization, they are related in that both tasks require taking an input article, then organizing and abstracting the information to generate a new document. Past work has developed methods for slide generation from papers Hu and Wan (2015); Li et al. (2021); Hu and Wan (2015); Fu et al. (2021), from code Wang et al. (2023a), or based on a query Sun et al. (2021). Poster generation has been explored in the form of content extraction for posters Xu and Wan (2021), interactive generation Wang et al. (2015), or full content generation using graphical models Qiang et al. (2016). To the best of our knowledge, our work is the first to create a unified method capable of generating a diverse range of templatic views of a source document.
Large Language Models (LLMs), which are central to our approach, have shown impressive capabilities in a variety of tasks Radford et al. (2019); Brown et al. (2020). Based on the transformer architecture Vaswani et al. (2017), LLMs have shown emergent abilities in tasks such as arithmetic and question answering Wei et al. (2022a). Similar to chain of thought prompting Wei et al. (2022b) and content planning prompting Wang et al. (2023b), we show that by generating an intermediate Structured Unified Representation (SURe) of an input document can improve performance over simply prompting the model to generate the final document Additionally, past work has explored using knowledge graphs to provide structure to LLMs in order to improve performance Pan et al. (2024); Rosset et al. (2020); Lewis et al. (2020). Similarly, we show that the structure provided by the SURe improves the quality of the final document, while being more easily adapted to new domains than a knowledge graph.
As past work has considered generation of templatic views to be separate tasks, methods for automatic evaluation of different templatic views has also been considered independently. LongSumm, the shared task introduced by Chandrasekaran et al. (2020), uses ROUGE scores to evaluate model performance Lin (2004). Fu et al. (2021) introduced Slide Level ROUGE to evaluate slide generation, which is a variant of ROUGE that contains a penalty for the number of slides. Qiang et al. (2016) used a trained regressor to predict the size of panels in a poster. In the summarization space, many automatic metrics for evaluation have been introduced such as BERTScore Zhang* et al. (2020), UniEval Zhong et al. (2022), BARTScore Yuan et al. (2021), BLANC Vasilyev et al. (2020), and MoverScore Zhao et al. (2019). However, these metrics are intended for a simple input document-summary setup, and do not take into account factors that affect the quality of other types of documents, such as structure and ordering. Our work is the first to introduce template agnostic evaluation, allowing uniform comparison of performance across template types.
Prompt Function Prompt
Generate the SURe
"Given the input text, extract the document title and authors.
For each section in the given input text, extract the most important sentences.
[…]
Output:"
[…]
"Summarize the following input in a <TEMPLATE TYPE>style.
Style parameters: <STYLE PARAMETERS>
[…]
Output:"
Table 1: Prompts used to generate the SURe and final LaTeX document. The template is pictured in Figure 2.
3 Data
Before introducing our approach to unified generation and evaluation of templatic document views, we begin by describing the data we use in this paper.
There is no existing dataset that includes multiple views of a single document. Instead, we evaluate our unified method, described in §4, on 3 datasets: DOC2PPT, LongSumm, and Paper-Poster Fu et al. (2021); Chandrasekaran et al. (2020); Qiang et al. (2016). These datasets are chosen because they each involve generating a different view of a document. We note that our method is not specific to the scientific domain. However, the scientific domain is one of the few domains with abundantly available public data of multiple templatic views of documents. We leave exploration of other domains to future work. The three datasets and their associated generation tasks are described below.
Slide Generation.
We use use the DOC2PPT dataset introduced by Fu et al. (2021). DOC2PPT is a dataset of 5.8K scientific papers in Computer Science and their respective slide decks. We randomly sample 1K examples from this dataset for evaluation. The slides are provided as an image for each slide. We use the Azure OCR tool to extract the text from each slide [^1] https://learn.microsoft.com/en-us/azure/ai-services/computer-vision/overview-ocr.
Blog Generation.
We use the LongSumm dataset introduced by Chandrasekaran et al. (2020), which includes blog posts of scientific papers in the Computer Science domain. LongSumm was originally introduced as a shared task for the purposes of long form summarization of scientific documents. Although LongSumm includes a blind set of 22 papers, our goal is to develop both a unified method of generation and evaluation across template types. As we are unable to compute our custom evaluation metric, described in §5, we exclude the blind test data from our evaluation. Instead, we use their released train split as our evaluation data, since our method does not require any training or supervision. Of the 531 released blog posts, we were able to access 505 of their papers. The other 26 papers included broken links or were behind a paywall. Longsumm also includes a set of extractive summaries from Lev et al. (2019). We exclude these summaries from our evaluation data.
Poster Generation.
We use the Paper-Poster dataset introduced by Qiang et al. (2016). Paper-Poster is a dataset of 85 papers in Computer Science and Biology, and their respective scientific posters. We exclude 2 examples containing corrupted PDFs from our evaluation. Although Qiang et al. (2016) uses the source files to extract the text of posters for evaluation, they only release the PDF formats. To preprocess the reference posters, we found that automatic tools to extract text from documents did not handle the visual layout of posters well, so we manually extracted the text of the posters in this dataset. Note that this process was only done to obtain metrics for evaluation, and that our unsupervised generation method is capable of creating target documents without the need for reference data, or this dataset-specific manual extraction step.
For all 3 datasets, we use the Azure Document Layout tool to extract the text of the input papers. [^2] https://learn.microsoft.com/en-us/azure/ai-services/document-intelligence/concept-layout?view=doc-intel-4.0.0
4 Unified LLM-powered Generation of Templatic Views
Our goal is to develop a method of generation that is: (1) unified, (2) adaptable, and (3) requires minimal supervision. We introduce a 2 step method: generate a Structured Unified Representation (SURe) and then generate the templatic view. By first generating a SURe, we are providing the model with a structured representation of the knowledge in the input document. This representation unifies the resulting templatic views, and can be used to generate any new template styles, achieving criteria (1) and (2). Additionally, this method is a zero shot method, meaning it requires no previously annotated data or training, achieving criteria (3). This process is visualized in Figure 1.
4.1 SURe Generation
The SURe is a JSON representation of the original document, consisting of a structured layout of the most important parts of the input. We provide the document to the model along with a template of the SURe. The prompt and template used can be found in Table 1 and Figure 2, respectively. By first generating a SURe, we are providing structure to the input document, while representing only the most important information, making it easier for the LLM to parse than the full input document. While we experiment with representing the SURe as a JSON in this paper, we note that a JSON format isn’t inherent to the SURe. The SURe could be represented in any format that provides structure to the input text (e.g. XML or Markdown) and we leave exploration of different formats to future work.
4.2 Templatic View Generation
We then feed the generated SURe as input back into the LLM, asking the model to generate the final output document. We ask the LLM to represent the final output document as a LaTeX file, as LaTeX is capable of formatting multiple styles of documents, including slides, posters, and blog posts. The prompts for each step can be found in Table 1. For each templatic view, the prompt to generate the final LaTeX document takes a short description of the desired output, which we refer to as style parameters. For the tasks we evaluate in this paper, we use the following style parameters:
  • Slides: “Slides should include a title page. Following slides should contain an informative slide title and short, concise bullet points. Longer slides should be broken up into multiple slides.”
  • Posters: “Posters should include a title section at the top. Each panel should include a heading and short, concise bullet points of the most important take-aways from that section.”
  • Blogs: “Blogs should include pargraphs introducing the topic, a summary of the input document, and important takeaways. The blog should be more readable to a general audience than the input document.”
By incorporating the use of style parameters, our method is adaptable to new templatic views; the user only needs to write a short description of the template style. This means that although we evaluate our method on the scientific domain, our method is not limited and could be generalized to other domains and template styles, such as enterprise-specific document types. We leave the exploration of other domains to future work.
Figure 2: Template provided to generate the SURe.
5 Template Agnostic Evaluation
As past work has treated the generation of different templatic views as separate tasks, the evaluation of these types has also been developed independently. Unfortunately, none of the previously used metrics scale well to other document types or more generally to a unified evaluation of templatic views. ROUGE Lin (2004); Chandrasekaran et al. (2020) is only designed to evaluate text against a text reference, and does not take structure and ordering of information into account; the same is true for other fuzzy text comparison metrics Zhang* et al. (2020); Zhong et al. (2022); Yuan et al. (2021). Slide-level ROUGE Fu et al. (2021) partly solves the problem with a length penalty but still does not account for distribution of information, nor is it easily extensible to formats other than slides. Finally using a trained regressor Qiang et al. (2016) as a metric is only possible with availability of human-labeled data, which is expensive and time-consuming to curate for new domains and applications.
Thus, our goal is to develop a framework of automatic evaluation that is template agnostic. In order to generalize to multiple templates, we introduce the concept of panels. A panel is a unit of organization within a document type, for which the placement and ordering of the panel is important to the overall flow of information in the document.
For example, we consider panels to be each slide in a slide deck and each section on a poster. We consider the entirety of a blog post to be a single panel. Although we test our method on the tasks of slide, blog, and poster generation, the concept of panels is not limited to these document types. For example, each post on a social media thread could be considered a panel, or each page on a website. By generalizing to the idea of panels, we can define an evaluation metric that is template agnostic.
We aim to unify the evaluation of templatic views by adapting and integrating prior metrics into a template agnostic precision-recall framework, which we refer to as Template-Agnostic Evaluation (TAE). TAE is not a new individual metric, but rather a framework of evaluation allowing generalization to new templates. For example, although ROUGE is not designed to capture posters, TAE can be used with ROUGE to evaluate poster generation.
The general formulation of TAE is as follows:
$\begin{split}\mathrm{Precision}=Q_{P}\times O_{P}\times L\\ \mathrm{Recall}=Q_{R}\times O_{R}\times L\end{split}$ (1)
in which $Q_{P}$ is the precision measure of quality (§5.1), $O_{P}$ is the precision penalty for order (§5.2), and $L$ is the penalty for length (§5.3). Similarly $Q_{R}$ is the recall measure of quality and $O_{R}$ is the recall penalty for order. The length penalty $L$ is non-reflexive, as discussed in §5.3. We calculate separate precision and recall scores, allowing evaluators to decide which measure is most important to them, including the ability to calculate an overall F-measure score.
5.1 Quality Measure
For the TAE precision score, we calculate the average similarity between the generated panels and their most similar reference panel. More specifically, we calculate $Q_{P}$ as follows:
$Q_{P}=\frac{1}{|\tilde{S}|}\sum_{\tilde{S}}\mathrm{max}_{\mathrm{sim}}(S,\tilde{S}_{i})$ (2)
in which $S$ is the set of reference panels and $\tilde{S}$ is the set of generated panels. For the similarity metric, the user can choose a similarity metric that best matches their use case, such as ROUGE, BERT-Score, or a custom trained regressor Lin (2004); Zhang* et al. (2020). For example, a user might choose ROUGE if they want a similarity metric that focuses on exact word overlap, or they may choose BERTScore to measure broader semantic similarity.
Similar to precision, for the TAE recall score, we calculate the average similarity between the reference panels and their most similar generated panel:
$Q_{R}=\frac{1}{|S|}\sum_{S}\mathrm{max}_{\mathrm{sim}}(\tilde{S},S_{i})$ (3)
In other words, we measure how similar each reference panel is to the generated panels.
By splitting the evaluation of quality into precision and recall, we can evaluate both the content of the slides that were generated as well as the coverage of this content against some reference.
5.2 Order Penalty
Broadly, the goal of the ordering penalty is to measure the similarity of the order of information in reference and generated panels, independent of other factors. Unfortunately, because the cardinality of panels in the two outputs is not necessarily the same, a direct one-to-one mapping to compare ordering is not feasible. Instead, a panel in one set can align to multiple references in the other, or none at all – as depicted in  Figure 3. Intuitively, our solution is to virtually replicate (resp. drop) panels that have multiple (resp. zero) alignments in the reference set so that a one-to-one mapping of ordering, can in fact be computed.
Formally, assume $S$ and $\tilde{S}$ are sequences of reference and generated panels respectively. We use the maximum similarity scores calculated in §5.1 to align the panels across sets. For the precision ordering penalty, we define the following operation $\lambda_{P}(s)=\sum_{\tilde{s}}\delta_{P}(s,\tilde{s})$, where
$\delta_{P}(s,\tilde{s})=\begin{cases}1,&\text{iff }s\rightarrow\tilde{s}\\ 0,&\text{otherwise}\end{cases}$
Intuitively, this captures the cardinality of the alignment of a given panel in $S$ with panels in $\tilde{S}$. Then, using this operation we can replace every panel $s\in S$ with $\lambda(s)$ copies, leading to an identical cardinality for both $S$ and $\tilde{S}$, and subsequent one-to-one mapping between their corresponding panels.
With this one-to-one mapping in place we can then measure the misalignments between the two sets of ordered panels to assign a penalty scores. To operationalize this, we associate panels with ranks in both sets and use a rank correlation metric to compute the degree of agreement. Specifically, rank assignment is done as follows: panels in $\tilde{S}$ are simply assigned ranks in order of appearance 1 through N – we call this $\tilde{S}_{ranking}^{P}$; meanwhile panels in $S$ are assigned the identical rank to their one-to-one aligned panel in $\tilde{S}$ and $\tilde{S}_{ranking}^{P}$ – we refer to these rankings as $S_{ranking}^{P}$. The final ordering penalty is computed using Spearman’s rank correlation Szmidt and Kacprzyk (2010):
$O_{P}=\frac{\mathrm{Spearman}(S_{ranking}^{P},\tilde{S}_{ranking}^{P})+1}{2}$ (4)
where we perform a linear transformation to map the original range of the correlation coefficient [-1, 1] to the desired range [0, 1] for precision.
Similarly, for the recall ordering penalty, we perform the same operation but in the other direction, mapping the reference panels to the generated panels. This means that panels st $\delta_{P}(s,\tilde{s})=0$ will be accounted for in the recall ordering penalty. Formally, $\lambda_{R}(s)=\sum_{s}\delta_{R}(\tilde{s},s)$, where
$\delta_{R}(\tilde{s},s)=\begin{cases}1,&\text{iff }\tilde{s}\rightarrow s\\ 0,&\text{otherwise}\end{cases}$
We calculate the rankings for recall similarly, assigning the panels in $S$ a rank of $S_{ranking}^{R}=[1,...,N]$ in order of appearance, while panels in $\tilde{S}$ are assigned the identical rank to their one-to-one aligned panel in $S$ and $S_{ranking}^{R}$. We then take the Spearman’s rank correlation as follows:
$O_{R}=\frac{\mathrm{Spearman}(S_{ranking}^{R},\tilde{S}_{ranking}^{R})+1}{2}$ (5)
An example of this process can be found in Figure 3.
Figure 3: Example of the process of obtaining the rankings for the precision ordering penalty. This process is reflexive, and panels not accounted for in the precision ordering penalty are accounted for in the recall ordering penalty.
Simularity Measure
Template Model Template R-1 R-L BERTScore
slides GPT4 SURe 8.95 8.37 38.08
Base 8.89 8.29 36.24
[…]
Base 5.62 5.09 31.64
[…]
Base 27.73 25.68 82.76
[…]
Base 27.02 25.10 82.85
[…]
Base 17.61 16.35 59.82
[…]
Base 8.38 7.55 35.45
Table 2: Evaluation results using GPT3.5 (gpt35-16k) and GPT4 (gpt4-32k). For each model, we compare the performance with and without the SURe. We report the TAE F1 scores as calculated in §5, using Rouge-1 (R-1), Rouge-L (R-L), and Bert-Score as the similarity metrics.
5.3 Length Penalty
Finally, we compute a length penalty for both the recall and precision scores. Similar to Fu et al. (2021), this is done as follows:
$L=e^{\frac{-\mathrm{abs}(|S|-|\tilde{S}|)}{|S|}}$ (6)
If $L$ were reflexive, the denominator in the reverse case would be $|\tilde{S}|$, meaning that as $|\tilde{S}|\rightarrow\infty$, $L\rightarrow 1$. In other words, the length penalty could be cheated by over-generating. Therefore, unlike the quality measure and ordering penalty, we chose not to make the length penalty reflexive.
6 Results
As a baseline, we skip the SURe generation step, passing the full text of the paper as input to the generate a LaTeX document prompt, rather than the SURe. We maintain the same style parameters for both settings. Past work on the Paper-Poster dataset and the DOC2PPT dataset did not release their code or results, so we cannot directly compare to their work. LongSumm’s test set is a blind set that does not allow us to compare using our custom metric.
We use gpt4-32k and gpt35-16k for our experiments. We truncated inputs that are too long for the input window and used a temperature of 0.0, based on hyper-parameter experiments described in §B.
6.1 Results of automatic evaluation
We report the results of our automatic evaluation in Table 2, computed the TAE scores as described in §5. As we can see from the results, first generating the SURe results in performance gains for most settings. We see that largest performance gains in the gpt35-16k experiments, indicating that our method has the largest benefits for simpler models. This is likely because more simple models have more difficulty parsing long unstructured documents, and the SURe provides a structured, filtered representation of the input document. The only setting we don’t see performance gains or similar performance is in the gpt4-32k setting on the poster template. Upon closer inspection, the references in this dataset are very verbose, averaging 391 tokens total. Our method produces generally less verbose posters, averaging 265 total tokens compared to 345 tokens produced by the baseline. We hypothesize that by editing the style parameters to include information about verbosity and length, we can improve performance on posters and leave exploration of this to future work.
6.1.1 Importance of Structure
Our method provides a template JSON to the LLM to generate the SURe, before passing the resulting JSON back to the LLM to generate the final document. A natural question is, how important is the JSON format? In other words, do we see performance gains as a result of the filtering of information, or is the structure provided by the JSON important as well? In order to test this question, we conduct an additional experiment in which we take the text within the SURe, discard the JSON structure, and pass the resulting body of text to the LLM to generate the final document. We randomly sample 100 documents each from LongSumm and Doc2PPT for the blog and poster generation tasks, respectively. We use the entirety of the Paper-Poster dataset, since it contains less than 100 examples. We use gpt35-16k for this experiment. The results of this experiment can be found in Table 3. As we can see from the results, passing the text of the SURe without the JSON structure decreases performance in most cases, with posters having the largest gap in performance, indicating that both the structure and content of the SURe are important for performance.
Template
SURe
Format
R-1 R-L
Bert
Score
slides JSON 7.7 7.3 35.1
Text 7.9 7.1 34.6
[…]
Text 28.3 24.4 82.5
[…]
Text 11.5 10.5 46.3
Table 3: TAE scores comparing the use of the JSON representation of the SURe or passing the text generated in the SURe, using gpt35-16k. We see that in most cases, the JSON representation performs better, indicating the importantance of structure.
Figure 4: Reasons annotators preferred each document. While annotators largely preferred documents generated with a SURe, the most common reasons for preference are better formatting and information content. We exclude the “Other” count as it was only selected once.
6.2 Human evaluation
We additionally conduct a human evaluation to test both our method and how well our evaluation framework aligns with human preference. We sample 100 documents each from Doc2PPT and LongSumm, and use the entirety of the Paper-Poster dataset. We present the annotators with 2 versions of each document, one generated with the SURe and one without. Both versions use gpt4-32k. The annotators are provided with the original paper and the intended document type (blog, slide deck, or poster), and are asked the following questions:
  1. 1.
    Which document do you prefer?
  2. 2.
    On a scale of 1-3, to what degree do you prefer your selection?
  3. 3.
    Why do you prefer your selection?
For Question 3, the annotators are given a multi-select checklist of reasons why they might prefer one document over the other. The checklist included (1) the quality of the content, (2) the formatting, (3) the document style matching the intended document types, (4) the information represented in the document, and (5) other. The “other” option included a free text box. The full annotation instructions, including a description of the reasons provided and examples, can be found in §A.
Each document was annotated by 3 different annotators. We employed 4 annotators from India, sourced via a third-party agency, to carry out the human evaluation of our task. A guideline document was provided to these evaluators, containing task-specific instructions, guidance, and several annotated examples. They were compensated at a rate of \$11.98 USD per hour for the total time that they spent working on the task, including a training round of annotation.
Which method do humans prefer?
The documents generated with a SURe were preferred by 82%, based on majority vote. 71% of the documents generated with a SURe were unanimously preferred by the annotators. We calculate a annotator agreement score of 0.51 with Krippendorff’s alpha, indicating that while this is a subjective and specialized task, even non-expert annotators agree to a moderate degree. A visualization of the reasons the annotators preferred their selection can be found in Figure 4. We can see from the results that while annotators largely preferred the documents generated with a SURe, the most common reasons for preference are better formatting and better information content. This indicates that the structure provided by the SURe makes it easier for the model to format the final document well. Additionally, the SURe only includes the most salient information from the original text, resulting in higher quality of information content. Finally, we see a fairly even distribution across the different templatic views for the reasons of preference, indicating humans prefer the documents generated with the SURe across different document types.
Metric PearsonR
TAE ROUGE-1 21.4
ROUGE-1 14.9
[…]
ROUGE-L 14.5
[…]
BERTScore 10.6*
Table 4: Correlation of evaluation metrics with human judgement. We compare each metric computed using the TAE framework versus the standard computation. *Indicates the correlation is not statistically significant ($p>0.01$)
Which metric correlates better with human judgement?
We use the human preference judgements to test if our metric, as described in §5, correlates more with human preference compared to prior evaluation metrics in the literature. For each annotation, given the degree of the preference $d$ we convert the model preference to a score $P$ as follows:
$P(d)=\begin{cases}3,&d=\textrm{strong}\\ 2,&d=\textrm{moderate}\\ 1,&d=\textrm{slight}\end{cases}$
If the annotator prefers the document generated without a SURe, we multiply $P(d)$ by $-1$. This allows us to measure if the metric captures both directionality of preference and degree.
Finally, we compute the automatic score for each metric, then calculate $S=m(\textrm{with SURe})-m(\textrm{skip SURe})$, where $m$ is the metric we are evaluating. If the annotator prefers the document generated without the SURe, we’d expect a good metric to assign a higher score to that document as well, resulting in both $S$ and $P$ being negative. Conversely, we’d expect instances where the opposite were true to result in positive scores for both $S$ and $P$. Using these intuitions we assign an affinity score of a metric with respect to human evaluation as the Pearson correlation Freedman et al. (2007) of $S$ and $P$.
Since prior metrics are not designed to account for the structure of the document, we compute them by extracting only the text of both the generated document and reference documents. The correlations with human judgement for each metric, computed standard compared to TAE variants can be found in Table 4. As we can see from the results, computing each evaluation metric using our template agnostic evaluation method correlates more highly with human judgement, except in the case of BERTScore. However, BERTScore is the only metric in which the correlation is not statistically significant, as it is likely simply a poor metric for this task. We hypothesize that BERTScore does not perform well on scientific data, since it is trained primarily on out of domain text.
7 Conclusion
In many domains, people choose to disseminate information across different modalities and formats, allowing for better communication of information to broader audiences. Past work has focused on developing methods to generate different templatic views of documents, tailoring the method to each template. In this paper, we introduced an LLM-powered method to unify the generation of templatic views of documents by first generating a Structured Unified Representation (SURe) of the input document, representing the underlying knowledge of the templatic views. We showed that our method results in improvements in performance, especially for a simpler model. We additionally introduced Template Agnostic Evaluation (TAE), a flexible precision-recall framework for automatic evaluation that is template agnostic, allowing for easy integration of existing evaluation metrics into a unified system. The framework is capable of jointly accounting for and measuring multiple facets of the output document including the quality of its content, its underlying structure and ordering, as well as its length. Finally, we conducted a human evaluation and showed that annotators prefer the documents generated using the SURe 82% of the time. We additionally showed that our framework of evaluation correlates more highly with human preference than standard evaluation metrics.
The use of templatic views of documents exists in many domains (e.g. creating a cover letter and personal website based on a resume), and in future work we will consider other domains of documents, as well as more complex scenarios, such as many-to-many document generation. Our work is a step towards a larger goal of generalized document generation.
Limitations
Although our methods are not domain specific, we only evaluated our methods in the scientific domain, due to the availability of public data. Further research is required to evaluate the performance of our methods in other domains. Additionally, our method is limited to textual content. Future work should explore incorporating multi-modal models and content generation. Finally, we evaluated our methods on closed source models, which may affect the reproducibility of our results.
Appendix A Annotation Instructions
A.1 Questions
Question 1 – Which document do you prefer? In this question, you are asked to choose which document version you prefer. Some examples of qualities you may use to decide your preference include:
  • The quality of the content – The text is grammatical and understandable. E.g. Document A contains major grammatical errors while Document B only contains minor errors.
  • The formatting – The formatting is reasonable and matches the formatting of the intended document type. E.g. A poster contains panels and each panel contains a header and body text.
  • The style – The document matches the style of the intended document type. E.g. Shorter, bulleted sentences in a slide deck.
  • Information represented in the document – The document contains sufficient information to represent the input document. E.g. A blog post represents the most important sections from the input document.
The above criteria are non-exhaustive. Not all criteria must be met, and you may use other relevant criteria to make your decision. You are not rating the document for factual correctness, and only need to refer to the corresponding scientific article if it will aid in making your preference. You can answer this question with either Document A or Document B.
Question 2 – On a scale of 1-3, to what degree do you prefer your selection?
In this question you will rate the degree to which you prefer your selection, on the following scale:
  1. 1.
    Small preference – The documents are similar in quality and only contain minor differences that affect my preference.
  2. 2.
    Moderate preference – I clearly prefer one document but the differences are not major.
  3. 3.
    Strong preference – I have a strong preference for one document and the differences between the documents are major.
Question 3 – Why do you prefer your selection? (You may select more than one property)
  • $\square$
    Formatting
  • $\square$
    Information
  • $\square$
    Quality
  • $\square$
    Style
  • $\square$
    Other (free text)
A.2 Edge cases
For most edge cases, it is up to your discretion on how to best handle the case. However, below are a few examples of how you could consider certain edge cases:
Example 1: Slides 1-5 of Document A are higher quality but slides 6-10 of Document B are higher quality. You could reason that the first slides represent the most important information, and choose Document A. However, since Document B contained higher quality slides for another portion of the document, you could rate your degree of preference as “Small preference.”
Example 2: Document A more closely matches the style of the intended document type, but Document B contains more relevant information to the source document. You could consider if Document A contains sufficient information to represent the input document, such as representing the most important sections. If yes, then you could prefer Document A. If not, then you could reason that information content is more important than style, and prefer Document B.
Example 3: Document A contains more relevant information than Document B, but also contains major formatting errors, such as text being cut off from the document.
You could reason that although Document A contains more relevant information, the major formatting errors are significant enough to prefer Document B.
Example 4: Neither document matches the style or formatting of the intended document type. Since neither document matches the style or formatting of the intended document type, you could consider other criteria, such as quality of content or the information represented.
Template Temp R-1 R-L BertScore
slides 0 7.7 7.3 35.1
0.25 7.8 7.2 35.4
0.5 7.6 7.0 36.4
0.75 7.5 7.0 35.5
1 8.0 7.4 35.8
[…]
0.25 27.2 25.5 82.6
0.5 28.0 26.2 82.8
0.75 27.2 25.4 82.6
1 26.6 24.8 82.7
[…]
0.25 14.2 13.0 53.1
0.5 13.6 12.5 52.2
0.75 13.2 12.0 50.8
1 12.6 11.6 50.3
Table 5: Results of the temperature hyperparameter experiments. There seems to be little consistency in performance as related to temperature across different templates.
Appendix B Temperature Experiments
We experiment with the temperature of the generations to see how temperature affects performance. Similar to the textual representation experiments, described in §6.1.1, we randomly sample 100 documents each from LongSumm and Doc2PPT for the blog and poster generation tasks, respectively. We use the entirety of the Paper-Poster dataset, since it contains less than 100 examples. We use gpt35-16k and experiment with the temperatures $[0.0,0.25,0.5,0.75,1.0]$. The results of this experiment can be found in Table 5. As we can see from the results, there seems to be little consistency across the different types in which temperature performs the best.

### v2

@@
 Abstract
-Authors seeking to communicate with broader audiences often compose their ideas about the same underlying knowledge in different documents and formats – for example, as slide decks, newsletters, reports, brochures, etc.
-Prior work in document generation has generally considered the creation of each separate format to be different a task, developing independent methods for generation and evaluation.
-This approach is suboptimal for the advancement of AI-supported content authoring from both research and application perspectives because it leads to fragmented learning processes, redundancy in models and methods, and disjointed evaluation.
-Thus, in our work, we consider each of these documents to be templatic views of the same underlying knowledge, and we aim to unify the generation and evaluation of these templatic views of documents.
-We begin by introducing an LLM-powered method to extract the most important information from an input document and represent this information in a structured format.
-We show that this unified representation can be used to generate multiple templatic views with no supervision and with very little guidance, improving over strong baselines.
-We additionally introduce a unified evaluation method that is template agnostic, and can be adapted to building document generators for heterogeneous downstream applications.
-Finally, we conduct a human evaluation, which shows that humans prefer 82% of the downstream documents generated with our method.
-Furthermore, the newly proposed evaluation metric correlates more highly with human judgement than prior metrics, while providing a unified evaluation method.
-Figure 1: Visualization of our method.
-Given an input document, we prompt the LLM to generate a Structure Unified Representation (SURe).
-We can use the SURe to prompt the model to generate a templatic view of the input document.
-We then evaluate the generations using our unified evaluation.
+Authors seeking to communicate with broader audiences often share their ideas in various document formats, such as slide decks, newsletters, reports, and posters.
+Prior work on document generation has generally tackled the creation of each separate format to be a different task, leading to fragmented learning processes, redundancy in models and methods, and disjointed evaluation.
+We consider each of these documents as templatic views of the same underlying knowledge/content, and we aim to unify the generation and evaluation of these templatic views.
+We begin by showing that current LLMs are capable of generating various document formats with little to no supervision.
+Further, a simple augmentation involving a structured intermediate representation can improve performance, especially for smaller models.
+We then introduce a novel unified evaluation framework that can be adapted to measuring the quality of document generators for heterogeneous downstream applications.
+This evaluation is adaptable to a range of user defined criteria and application scenarios, obviating the need for task specific evaluation metrics.
+Finally, we conduct a human evaluation, which shows that people prefer 82% of the documents generated with our method, while correlating more highly with our unified evaluation framework than prior metrics in the literature.
+Figure 1: Visualization of our method to unify the generation and evaluation of templatic views of documents.
+Given an input document, we prompt the LLM to generate an intermediate representation.
+We can use the representation to prompt the model to generate a templatic view of the input document.
+We then evaluate the generations using our unified evaluation framework.
+The LLM represented in the figure is the same model.
 1 Introduction
-Sharing information in multiple formats is an important part of communication in a variety of domains, as it allows for knowledge to be disseminated to a wider audience.
-A product manager may need to create a requirements document, a product pitch deck, and a product announcement newsletter.
+Sharing information is vital for communication and discourse across domains, as it allows for knowledge to be disseminated to a wider audience.
+This is often done by users through documents in multiple formats that nevertheless share some underlying knowledge.
+A product manager may need to create a requirements spec, a product pitch deck, and an announcement newsletter for the same project.
 Likewise, a person on the job market may create a resume, a cover letter, and a personal website.
 We consider these documents to be templatic views of the same underlying knowledge.
-This is also true for the scientific domain, in which researchers need to create documents in multiple formats to effectively disseminate their work, – such as through conference talks, social media posts, poster presentations, and non-technical blog posts.
-Sharing knowledge in multiple formats broadens its audience and can help bridge the information gap between domain experts, researchers in adjacent fields, and even the general public – leading to greater understanding, new collaboration and accelerated progress in the field.
-With the significant increase in the pace of publication in recent years Bornmann and Mutz (2014), it has become essential for scientists to engage with the community and share their work in more compact, efficient, or easier to digest formats.
-Past work on document generation has focused on developing methods of generation and evaluation for each specific document type individually Fu et al.
+This is equally true for the scientific domain, in which researchers create documents in multiple formats to effectively communicate and showcase their work, – such as through academic papers, conference talks, social media posts, poster presentations, and non-technical blog posts.
+Sharing knowledge in multiple formats broadens the audience and can help bridge the information gap between domain experts, researchers in adjacent fields, and even the general public, leading to greater understanding, collaborations and accelerated progress Bornmann and Mutz (2014).
+Past work on document generation has focused on developing generation and evaluation methods specific to a single document type Fu et al.
 (2021); Qiang et al.
@@
 (2020).
-However, tailoring methods to each template is time consuming and limiting, resulting in methods that are usable only for each intended template that cannot be compared across template types with specialized evaluation metrics.
-Our goal is to unify the tasks of generating and evaluating templatic views of documents.
-By doing so, we aim to allow system designers to adapt to new document types, templates and domains easily and efficiently, and to allow researchers to compare approaches more flexibly and fairly.
-We begin by setting a few desiderata for our system for templatic view generation; it must be (1) unified, (2) adaptable, and (3) require minimal supervision.
-In order to design such a system to generate each of these templates, the method must be able to create structure and follow style norms associated with each template.
-We thus introduce an LLM-powered approach to generate a Structure Unified Representation (SURe) that represents the most important knowledge of an input document in a structured format.
-We demonstrate that this SURe can be used with no supervision and with very little prompted guidance about the output formats, to generate multiple templatic views of a document.
-Meanwhile, automatic evaluation is crucial for researchers and developers to iterate more quickly than they otherwise would be capable of by relying solely on human evaluation.
-However, because in the past each templatic view has been tackled independently, previous work has used different evaluation metrics for each task, such as BERTScore for summarization Zhang* et al.
-(2020), Slide Level ROUGE for slide decks Qiang et al.
-(2016), or even a custom-trained, model-based metric for posters Wang et al.
+Narrow, custom methods tailored to individual document types are, nevertheless, time consuming to engineer and manage over the long term.
+For example, in an enterprise setting, it’s common to have dozens of occupation- and task-specific documents, each with their own template.Additionally, specific trained methods require data that may be expensive to acquire, or even be unavailable entirely.
+Meanwhile, LLMs have recently shown great success in long document generation Radford et al.
+(2019); Brown et al.
+(2020), indicating that this fragmentation of methods may no longer be necessary.
+Thus, our goal is to unify methods for both generating and evaluating templatic views of documents, allowing system designers and engineers to manage and adapt to a range of document types and domains easily and efficiently.
+We begin by showing that LLMs are capable of diverse, structured document generation, requiring very little instructional guidance to do so effectively.
+Additionally, a few minor augmentations to the prompt – such as a structured, intermediate representation, and simple stylistic descriptions – can further improve downstream performance, especially for smaller, less resource intensive models.
+These findings have important implications on the deployment and scaling of unified, real-world AI-assisted document authoring systems.
+In similar vein, we then introduce Template Adaptable Evaluation (TAE), departing from prior work’s task specific evaluation methods Zhang* et al.
+(2020); Qiang et al.
+(2016); Wang et al.
 (2015).
-In this paper, we introduce Template Agnostic Evaluation (TAE) – a unified precision-recall style method of automatic evaluation that is template agnostic.
-Importantly, this evaluation framework is highly customizable, allowing users to easily integrate existing text-based metrics from the literature into its formulation and tailor its function to their specific use case.
-We evaluate our unified approach to templatic view generation and evaluation on 3 types of documents: slides, posters, and blog posts.
-In each of these, the inputs are scientific articles and the outputs are expected to be documents of the corresponding types.
-We use the DOC2PPT Fu et al.
-(2021), Paper-Poster Qiang et al.
-(2016), and LongSumm Chandrasekaran et al.
-(2020) datasets, respectively.
-We show that our method of generating a Structured Unified Representation leads to improvements in performance across each tasks, demonstrating that LLMs are capable of generating templatic views with structure-augmented prompts, despite having no supervision.
-We additionally conduct a human evaluation to validate both our document generation method and evaluation metric.
-We show that humans prefer the output yielded by our method 82% of the time and that our evaluation metric correlates more highly with human preference than other popular metrics, such as ROUGE and BERTScore Lin (2004); Zhang* et al.
+TAE is a unified precision-recall style framework for automatic evaluation that is highly customizable, allowing users to easily integrate existing text-based metrics from the literature into its formulation and tailor it to their specific use case.Additionally, this framework allows developers to compare performance across document types, without needing to develop an evaluation metric for each individual template.
+We evaluate our unified approach for templatic view generation and evaluation on 3 types of documents: slides, posters, and blog posts Fu et al.
+(2021); Qiang et al.
+(2016); Chandrasekaran et al.
 (2020).
-While the focus of this paper is the transformation of scientific documents, this choice was primarily driven by the availability of publicly available data.
-We note that nothing about our unified framework is domain specific, and we believe our approach for the creation and evaluation of templatic views of documents is easily adaptable to novel domains, and document types.
-More generally, we believe our work is a step towards the broader and more challenging goal of AI-assisted authoring support, and future research in reasoning over and generation of long documents.
+Our experiments demonstrate that using a structured intermediate representation leads to improvements in performance across tasks, with greater gains for smaller language language models.
+In our human evaluation to validate both our unified document generation method and evaluation metric, we show that annotators prefer the output yielded by the structure-aware generation process 82% of the time and that our evaluation metric correlates more highly with human preference than other popular metrics.
+We release our code [^1] https://github.com/microsoft/knowledge-centric-templatic-views to support future research.
 2 Related Work
@@
 (2022b) and content planning prompting Wang et al.
-(2023b), we show that by generating an intermediate Structured Unified Representation (SURe) of an input document can improve performance over simply prompting the model to generate the final document Additionally, past work has explored using knowledge graphs to provide structure to LLMs in order to improve performance Pan et al.
-(2024); Rosset et al.
-(2020); Lewis et al.
-(2020).
-Similarly, we show that the structure provided by the SURe improves the quality of the final document, while being more easily adapted to new domains than a knowledge graph.
-As past work has considered generation of templatic views to be separate tasks, methods for automatic evaluation of different templatic views has also been considered independently.
+(2023b), we show that by generating an intermediate representation of an input document can improve performance over simply prompting the model to generate the final document from the original input.
+As past work has tackled generation of templatic views as separate tasks, methods for automatic evaluation of different document types is fragmented.
 LongSumm, the shared task introduced by Chandrasekaran et al.
-(2020), uses ROUGE scores to evaluate model performance Lin (2004).
+(2020), uses ROUGE to evaluate model performance Lin (2004).
 Fu et al.
-(2021) introduced Slide Level ROUGE to evaluate slide generation, which is a variant of ROUGE that contains a penalty for the number of slides.
+(2021) introduced Slide Level ROUGE to evaluate slide generation, a variant that contains a penalty for the number of slides.
 Qiang et al.
-(2016) used a trained regressor to predict the size of panels in a poster.
-In the summarization space, many automatic metrics for evaluation have been introduced such as BERTScore Zhang* et al.
+(2016) used a trained regressor.
+For summarization, many automatic evaluation metrics have been introduced such as BERTScore Zhang* et al.
 (2020), UniEval Zhong et al.
@@
 (2019).
-However, these metrics are intended for a simple input document-summary setup, and do not take into account factors that affect the quality of other types of documents, such as structure and ordering.
-Our work is the first to introduce template agnostic evaluation, allowing uniform comparison of performance across template types.
-Prompt Function Prompt
-Generate the SURe
-"Given the input text, extract the document title and authors.
-For each section in the given input text, extract the most important sentences.
-[…]
-[…]
-[…]
-Output:"
-[…]
-"Summarize the following input in a <TEMPLATE TYPE>style.
-Style parameters: <STYLE PARAMETERS>
-[…]
-[…]
-Output:"
-Table 1: Prompts used to generate the SURe and final LaTeX document. The template is pictured in Figure 2.
+However, these metrics are intended for a simple input document-summary setup, and do not take into account factors that affect the quality of other types of documents (e.g.
+structure).
+Our work is the first to introduce template adaptable evaluation, allowing uniform comparison of performance across template types.
 3 Data
-Before introducing our approach to unified generation and evaluation of templatic document views, we begin by describing the data we use in this paper.
+We begin by describing the data used in this paper.
 There is no existing dataset that includes multiple views of a single document.
-Instead, we evaluate our unified method, described in §4, on 3 datasets: DOC2PPT, LongSumm, and Paper-Poster Fu et al.
+Instead, we evaluate our unified method, described in §4, on 3 existing datasets: DOC2PPT, LongSumm, and Paper-Poster Fu et al.
 (2021); Chandrasekaran et al.
@@
 These datasets are chosen because they each involve generating a different view of a document.
-We note that our method is not specific to the scientific domain.
-However, the scientific domain is one of the few domains with abundantly available public data of multiple templatic views of documents.
-We leave exploration of other domains to future work.
+Although our method is not specific to the scientific domain, it is one of the few domains with abundantly available public data of multiple templatic views  [^2] We acknowledge that scientific writing does have structural regularities that may influence unified document generation.
+Due to the lack of other available datasets we leave exploration of other domains to future work..
 The three datasets and their associated generation tasks are described below.
 Slide Generation.
-We use use the DOC2PPT dataset introduced by Fu et al.
-(2021).
-DOC2PPT is a dataset of 5.8K scientific papers in Computer Science and their respective slide decks.
-We randomly sample 1K examples from this dataset for evaluation.
+We use use the DOC2PPT dataset Fu et al.
+(2021), which contains 5.8K scientific papers in Computer Science and their respective slide decks.
+As Fu et al.
+(2021) do not release data splits or code, we randomly sample 1K examples from this dataset for evaluation.
 The slides are provided as an image for each slide.
@@
 Blog Generation.
-We use the LongSumm dataset introduced by Chandrasekaran et al.
+We use the LongSumm dataset Chandrasekaran et al.
 (2020), which includes blog posts of scientific papers in the Computer Science domain.
-LongSumm was originally introduced as a shared task for the purposes of long form summarization of scientific documents.
-Although LongSumm includes a blind set of 22 papers, our goal is to develop both a unified method of generation and evaluation across template types.
-As we are unable to compute our custom evaluation metric, described in §5, we exclude the blind test data from our evaluation.
-Instead, we use their released train split as our evaluation data, since our method does not require any training or supervision.
-Of the 531 released blog posts, we were able to access 505 of their papers.
-The other 26 papers included broken links or were behind a paywall.
-Longsumm also includes a set of extractive summaries from Lev et al.
-(2019).
-We exclude these summaries from our evaluation data.
+Since our approach requires no training or supervision, we use the entire training split from Longsumm as our evaluation set.
+Of the 531 publicly released blog posts in this set, we could only access 505, with the other 26 including broken links or being behind a paywall.
+Notably, while Longsumm includes a blind test set of 22 papers, this test set only consists of inputs without their reference outputs, thus making it impossible to compute our custom evaluation metric (see §5).
+In the interest of completeness and comparison to prior work, we do, however submit runs from our systems to the leader board and report the results of this blind test set in Appendix D.
 Poster Generation.
-We use the Paper-Poster dataset introduced by Qiang et al.
-(2016).
-Paper-Poster is a dataset of 85 papers in Computer Science and Biology, and their respective scientific posters.
-We exclude 2 examples containing corrupted PDFs from our evaluation.
+We use the Paper-Poster dataset Qiang et al.
+(2016), which consists of a dataset of 85 papers in Computer Science and Biology, and their respective scientific posters; two examples containing corrupted PDFs are excluded.
 Although Qiang et al.
-(2016) uses the source files to extract the text of posters for evaluation, they only release the PDF formats.
+(2016) release data splits, they do not release code or results for comparison.
+Given the small size of the dataset, we use it in its entirety for more robust results.
+While the authors uses the source files to extract the text of posters for evaluation, they only release the PDF formats.
 To preprocess the reference posters, we found that automatic tools to extract text from documents did not handle the visual layout of posters well, so we manually extracted the text of the posters in this dataset.
-Note that this process was only done to obtain metrics for evaluation, and that our unsupervised generation method is capable of creating target documents without the need for reference data, or this dataset-specific manual extraction step.
-For all 3 datasets, we use the Azure Document Layout tool to extract the text of the input papers.
-[^2] https://learn.microsoft.com/en-us/azure/ai-services/document-intelligence/concept-layout?view=doc-intel-4.0.0
+Note that this process was only done to obtain evaluation scores, and that our unsupervised generation method is capable of creating target documents without the need for reference data.
+For all 3 datasets, we use the Azure Document Layout tool to extract the text of the input papers. [^4] https://learn.microsoft.com/en-us/azure/ai-services/document-intelligence/concept-layout
 4 Unified LLM-powered Generation of Templatic Views
-Our goal is to develop a method of generation that is: (1) unified, (2) adaptable, and (3) requires minimal supervision.
-We introduce a 2 step method: generate a Structured Unified Representation (SURe) and then generate the templatic view.
-By first generating a SURe, we are providing the model with a structured representation of the knowledge in the input document.
-This representation unifies the resulting templatic views, and can be used to generate any new template styles, achieving criteria (1) and (2).
-Additionally, this method is a zero shot method, meaning it requires no previously annotated data or training, achieving criteria (3).
-This process is visualized in Figure 1.
-4.1 SURe Generation
-The SURe is a JSON representation of the original document, consisting of a structured layout of the most important parts of the input.
-We provide the document to the model along with a template of the SURe.
-The prompt and template used can be found in Table 1 and Figure 2, respectively.
-By first generating a SURe, we are providing structure to the input document, while representing only the most important information, making it easier for the LLM to parse than the full input document.
-While we experiment with representing the SURe as a JSON in this paper, we note that a JSON format isn’t inherent to the SURe.
-The SURe could be represented in any format that provides structure to the input text (e.g.
-XML or Markdown) and we leave exploration of different formats to future work.
-4.2 Templatic View Generation
-We then feed the generated SURe as input back into the LLM, asking the model to generate the final output document.
-We ask the LLM to represent the final output document as a LaTeX file, as LaTeX is capable of formatting multiple styles of documents, including slides, posters, and blog posts.
-The prompts for each step can be found in Table 1.
-For each templatic view, the prompt to generate the final LaTeX document takes a short description of the desired output, which we refer to as style parameters.
-For the tasks we evaluate in this paper, we use the following style parameters:
-  • Slides: “Slides should include a title page. Following slides should contain an informative slide title and short, concise bullet points. Longer slides should be broken up into multiple slides.”
-  • Posters: “Posters should include a title section at the top. Each panel should include a heading and short, concise bullet points of the most important take-aways from that section.”
-  • Blogs: “Blogs should include pargraphs introducing the topic, a summary of the input document, and important takeaways.
-The blog should be more readable to a general audience than the input document.”
-By incorporating the use of style parameters, our method is adaptable to new templatic views; the user only needs to write a short description of the template style.
-This means that although we evaluate our method on the scientific domain, our method is not limited and could be generalized to other domains and template styles, such as enterprise-specific document types.
-We leave the exploration of other domains to future work.
-Figure 2: Template provided to generate the SURe.
-5 Template Agnostic Evaluation
-As past work has treated the generation of different templatic views as separate tasks, the evaluation of these types has also been developed independently.
-Unfortunately, none of the previously used metrics scale well to other document types or more generally to a unified evaluation of templatic views.
-ROUGE Lin (2004); Chandrasekaran et al.
-(2020) is only designed to evaluate text against a text reference, and does not take structure and ordering of information into account; the same is true for other fuzzy text comparison metrics Zhang* et al.
-(2020); Zhong et al.
-(2022); Yuan et al.
-(2021).
-Slide-level ROUGE Fu et al.
-(2021) partly solves the problem with a length penalty but still does not account for distribution of information, nor is it easily extensible to formats other than slides.
-Finally using a trained regressor Qiang et al.
-(2016) as a metric is only possible with availability of human-labeled data, which is expensive and time-consuming to curate for new domains and applications.
-Thus, our goal is to develop a framework of automatic evaluation that is template agnostic.
+The most straightforward way to transform documents between templatic views using LLMs, is to simply prompt the system to generate the target view given the input.
+However, similar to chain of thought prompting Wei et al.
+(2022b), we hypothesize that first generating a structured, intermediate representation of an input document and then reasoning over that representation will result in better generations than directly prompting the model.
+Our goal is to evaluate the capabilites of LLMs to generate long, structured documents, and experiment with how structured prompting can improve performance.
+We experiment with a simple general two-step process: first generate an intermediate representation, then generate the templatic view.
+These steps are described in greater detail below, and the process is visualized in Figure 1.
+Intermediate Representation Generation.
+In this work, we set the intermediate representation to be a JSON consisting of a structured layout of the most important parts of the input.
+We provide the input document to the model along with a template of the representation and prompt it to extract the most important information from the input document, and format it in the given JSON structure.
+The exact prompts and JSON structure can be found in Appendix §A.
+While our experiments use a JSON intermediate representation, note that other formats that provide structure to the input text could be employed (e.g.
+XML or Markdown).
+Rather than trying to optimize for the best representation format, our goal is to show that this chain of extraction approach along with structured augmentation to prompts can aid the quality of generations from LLMs.
+We leave exploration of different formats and other prompt optimization to future work.
+Templatic View Generation.
+We then feed the generated representation as input back into the LLM, prompting the model to generate the final output document, represented as a LaTeX document.
+For each templatic view, the prompt to generate the final LaTeX document takes a short description of the desired output, which we refer to as a style parameter.
+For example, the style parameter for slide generation is as follows: “Slides should include a title page.
+Following slides should contain an informative slide title and short, concise bullet points.
+Longer slides should be broken up into multiple slides.” The use of style parameters makes our method adaptable to new templatic views; the user only needs to write a short description of the template style.
+Both the generation of the intermediate representations and the final documents require little to no prompt engineering.
+The prompts and style parameters can be found in Appendix §A.
+5 Template Adaptable Evaluation
+Prior work on document generation has treated the evaluation of different templatic views as separate tasks.
+Thus, our goal is to develop a framework of automatic evaluation that is template adaptable.
+This not only allows us to compare performance across diverse datasets, it also removes the requirement of designing and maintaining individual metrics for each template.
 In order to generalize to multiple templates, we introduce the concept of panels.
@@
 For example, each post on a social media thread could be considered a panel, or each page on a website.
-By generalizing to the idea of panels, we can define an evaluation metric that is template agnostic.
-We aim to unify the evaluation of templatic views by adapting and integrating prior metrics into a template agnostic precision-recall framework, which we refer to as Template-Agnostic Evaluation (TAE).
-TAE is not a new individual metric, but rather a framework of evaluation allowing generalization to new templates.
-For example, although ROUGE is not designed to capture posters, TAE can be used with ROUGE to evaluate poster generation.
-The general formulation of TAE is as follows:
+We aim to unify the evaluation of templatic views by integrating prior metrics into a template adaptable precision-recall framework, which we refer to as Template-Adaptable Evaluation (TAE).
+TAE is not a new individual metric, but rather an evaluation framework that allows generalization to new templates.
+For example, TAE can even be used with ROUGE to evaluate poster generation.
+The general TAE formulation is as follows:
 $\begin{split}\mathrm{Precision}=Q_{P}\times O_{P}\times L\\ \mathrm{Recall}=Q_{R}\times O_{R}\times L\end{split}$ (1)
-in which $Q_{P}$ is the precision measure of quality (§5.1), $O_{P}$ is the precision penalty for order (§5.2), and $L$ is the penalty for length (§5.3).
-Similarly $Q_{R}$ is the recall measure of quality and $O_{R}$ is the recall penalty for order.
-The length penalty $L$ is non-reflexive, as discussed in §5.3.
-We calculate separate precision and recall scores, allowing evaluators to decide which measure is most important to them, including the ability to calculate an overall F-measure score.
+in which $Q_{P}$ is the precision measure of quality (§5.1), $O_{P}$ is the precision penalty for order (§5.2), and $L$ is the non-reflexive penalty for length (§5.3).
+Similarly, $Q_{R}$ is the recall quality measure and $O_{R}$ is the recall penalty for order.
+The precision-recall formulation allows evaluators to decide which measure is most important to them, or calculate an overall F-measure score.
 5.1 Quality Measure
-For the TAE precision score, we calculate the average similarity between the generated panels and their most similar reference panel. More specifically, we calculate $Q_{P}$ as follows:
+For the TAE precision score, we calculate the average similarity between the generated panels and their most similar reference panel as follows:
 $Q_{P}=\frac{1}{|\tilde{S}|}\sum_{\tilde{S}}\mathrm{max}_{\mathrm{sim}}(S,\tilde{S}_{i})$ (2)
 in which $S$ is the set of reference panels and $\tilde{S}$ is the set of generated panels.
-For the similarity metric, the user can choose a similarity metric that best matches their use case, such as ROUGE, BERT-Score, or a custom trained regressor Lin (2004); Zhang* et al.
+For the similarity metric, the user can choose a metric that best matches their use case, such as ROUGE, BERT-Score, or a custom trained regressor Lin (2004); Zhang* et al.
 (2020).
-For example, a user might choose ROUGE if they want a similarity metric that focuses on exact word overlap, or they may choose BERTScore to measure broader semantic similarity.
-Similar to precision, for the TAE recall score, we calculate the average similarity between the reference panels and their most similar generated panel:
+For example, a user might choose ROUGE if they want a similarity metric that focuses on exact word overlap, or BERTScore to measure broader semantic similarity.
+Similar to precision, the TAE recall score is calculated as the average similarity between the reference panels and their most similar generated panel:
 $Q_{R}=\frac{1}{|S|}\sum_{S}\mathrm{max}_{\mathrm{sim}}(\tilde{S},S_{i})$ (3)
-In other words, we measure how similar each reference panel is to the generated panels.
 By splitting the evaluation of quality into precision and recall, we can evaluate both the content of the slides that were generated as well as the coverage of this content against some reference.
@@
 zero) alignments in the reference set so that a one-to-one mapping of ordering, can in fact be computed.
-Formally, assume $S$ and $\tilde{S}$ are sequences of reference and generated panels respectively.
-We use the maximum similarity scores calculated in §5.1 to align the panels across sets.
+Formally, assume $S$ and $\tilde{S}$ are sequences of reference and generated panels respectively. We use the maximum similarity scores calculated in §5.1 to align the panels across sets.
 For the precision ordering penalty, we define the following operation $\lambda_{P}(s)=\sum_{\tilde{s}}\delta_{P}(s,\tilde{s})$, where
 $\delta_{P}(s,\tilde{s})=\begin{cases}1,&\text{iff }s\rightarrow\tilde{s}\\ 0,&\text{otherwise}\end{cases}$
-Intuitively, this captures the cardinality of the alignment of a given panel in $S$ with panels in $\tilde{S}$.
-Then, using this operation we can replace every panel $s\in S$ with $\lambda(s)$ copies, leading to an identical cardinality for both $S$ and $\tilde{S}$, and subsequent one-to-one mapping between their corresponding panels.
-With this one-to-one mapping in place we can then measure the misalignments between the two sets of ordered panels to assign a penalty scores.
-To operationalize this, we associate panels with ranks in both sets and use a rank correlation metric to compute the degree of agreement.
+Intuitively, this captures the cardinality of the alignment of a panel in $S$ with panels in $\tilde{S}$.
+Then, using this operation we can replace every $s\in S$ with $\lambda_{P}(s)$ copies, leading to an identical cardinality for both $S$ and $\tilde{S}$, and subsequent one-to-one mapping between their corresponding panels.
+Figure 2: Example of the process of obtaining the rankings for the precision ordering penalty.
+We first use the similarity measure to map each generated panel to its most similar reference document.
+This mapping is used to calculate the precision quality score $Q_{P}$.
+We then use the mappings to create a one-to-one alignment from the generated to the reference panels, which we use to calculate the precision ordering penalty ($O_{P}$).
+By creating a one-to-one alignment, we are able to represent inversions in the ordering.
+This process is reflexive, and panels not accounted for in the precision ordering penalty are accounted for in the recall ordering penalty.
+Then, to operationalize a penalty score for the two sets of ordered panels we associate them with ranks in both sets and use a rank correlation metric to compute the degree of agreement.
 Specifically, rank assignment is done as follows: panels in $\tilde{S}$ are simply assigned ranks in order of appearance 1 through N – we call this $\tilde{S}_{ranking}^{P}$; meanwhile panels in $S$ are assigned the identical rank to their one-to-one aligned panel in $\tilde{S}$ and $\tilde{S}_{ranking}^{P}$ – we refer to these rankings as $S_{ranking}^{P}$.
+An example of this process can be found in Figure 2.
 The final ordering penalty is computed using Spearman’s rank correlation Szmidt and Kacprzyk (2010):
 $O_{P}=\frac{\mathrm{Spearman}(S_{ranking}^{P},\tilde{S}_{ranking}^{P})+1}{2}$ (4)
-where we perform a linear transformation to map the original range of the correlation coefficient [-1, 1] to the desired range [0, 1] for precision.
-Similarly, for the recall ordering penalty, we perform the same operation but in the other direction, mapping the reference panels to the generated panels.
-This means that panels st $\delta_{P}(s,\tilde{s})=0$ will be accounted for in the recall ordering penalty.
-Formally, $\lambda_{R}(s)=\sum_{s}\delta_{R}(\tilde{s},s)$, where
-$\delta_{R}(\tilde{s},s)=\begin{cases}1,&\text{iff }\tilde{s}\rightarrow s\\ 0,&\text{otherwise}\end{cases}$
-We calculate the rankings for recall similarly, assigning the panels in $S$ a rank of $S_{ranking}^{R}=[1,...,N]$ in order of appearance, while panels in $\tilde{S}$ are assigned the identical rank to their one-to-one aligned panel in $S$ and $S_{ranking}^{R}$.
-We then take the Spearman’s rank correlation as follows:
-$O_{R}=\frac{\mathrm{Spearman}(S_{ranking}^{R},\tilde{S}_{ranking}^{R})+1}{2}$ (5)
-An example of this process can be found in Figure 3.
-Figure 3: Example of the process of obtaining the rankings for the precision ordering penalty.
-This process is reflexive, and panels not accounted for in the precision ordering penalty are accounted for in the recall ordering penalty.
-Simularity Measure
-Template Model Template R-1 R-L BERTScore
-slides GPT4 SURe 8.95 8.37 38.08
-Base 8.89 8.29 36.24
-[…]
-Base 5.62 5.09 31.64
-[…]
-Base 27.73 25.68 82.76
-[…]
-Base 27.02 25.10 82.85
-[…]
-Base 17.61 16.35 59.82
-[…]
-Base 8.38 7.55 35.45
-Table 2: Evaluation results using GPT3.5 (gpt35-16k) and GPT4 (gpt4-32k).
-For each model, we compare the performance with and without the SURe.
-We report the TAE F1 scores as calculated in §5, using Rouge-1 (R-1), Rouge-L (R-L), and Bert-Score as the similarity metrics.
+where we perform a linear transformation to map the original range of the correlation coefficient [-1, 1] to the desired range [0, 1].
+Similarly, for the recall ordering penalty, we map the reference panels to the generated panels, calculated as $\lambda_{R}(s)=\sum_{s}\delta_{R}(\tilde{s},s)$.
+$O_{R}$ is calculated similar to $O_{P}$, using the recall rankings.
 5.3 Length Penalty
@@
 $L=e^{\frac{-\mathrm{abs}(|S|-|\tilde{S}|)}{|S|}}$ (5)
-If $L$ were reflexive, the denominator in the reverse case would be $|\tilde{S}|$, meaning that as $|\tilde{S}|\rightarrow\infty$, $L\rightarrow 1$.
-In other words, the length penalty could be cheated by over-generating.
-Therefore, unlike the quality measure and ordering penalty, we chose not to make the length penalty reflexive.
+We chose to keep $L$ non-reflexive, because in the reverse case – as $|\tilde{S}|\rightarrow\infty$, $L\rightarrow 1$ – the metric could be cheated by over-generating.
 6 Results
-As a baseline, we skip the SURe generation step, passing the full text of the paper as input to the generate a LaTeX document prompt, rather than the SURe.
-We maintain the same style parameters for both settings.
-Past work on the Paper-Poster dataset and the DOC2PPT dataset did not release their code or results, so we cannot directly compare to their work.
-LongSumm’s test set is a blind set that does not allow us to compare using our custom metric.
-We use gpt4-32k and gpt35-16k for our experiments. We truncated inputs that are too long for the input window and used a temperature of 0.0, based on hyper-parameter experiments described in §B.
+As mentioned in §3, past work on Doc2PPT and Paper-Posters do not release code, making it difficult to do a direct comparison.
+They also do not report any baselines to compare against.
+Meanwhile, Longsumm’s blind test does not allow us to compute our custom metric, although we do report the leaderboard results in Appendix D.
+Notably, with almost no prompt engineering our LLM-based system places second on this leaderboard.
+We argue that for the investigation in this paper, direct comparison to prior non-LLM baselines is not only unfair to those approaches, but not particularly insightful.
+Therefore, similar to Wei et al.
+(2022b), we focus on variants of our LLM-based method and treat them as baselines.
+Example outputs of each template type can be found in Appendix E.
+We conduct experiments with the following settings: (1) No Representation – this is the default setting of going directly from the source document to the target document.
+We skip the intermediate generation step, passing the full paper as input.
+We experiment both with and without the style parameters.
+(2) Own Representation – we do not pass a JSON structure to the intermediate generation step, and allow the model to choose its own structure.
+(3) Text Representation – we extract the text from the intermediate representation, discarding the JSON structure.
+(4) JSON Representation – this is the full JSON structure for the intermediate generation step.
+We experiment both with and without the style parameters.
+We use gpt35-16k in our main set of experiments.
+We truncate text that is too long for the input window and use a temperature of 0.0 as standard.
+[^5] A detailed evaluation of the temperature hyper-parameter is included in Appendix §C
+Similarity Measure
+Rep. Style R-L M B BERTS
+Slides None $\times$ 5.0 6.4 0.3 31.6
+None $\checkmark$ 5.1 6.0 0.4 31.7
+Own $\checkmark$ 6.5 7.1 1.2 36.1
+Text $\checkmark$ 7.3 8.0 1.4 36.4
+JSON $\times$ 4.2 6.0 0.3 31.4
+JSON $\checkmark$ 7.4 8.4 1.5 36.9
+[…]
+None $\checkmark$ 25.1 17.7 2.3 82.8
+Own $\checkmark$ 23.9 19.2 2.3 82.2
+Text $\checkmark$ 25.4 19.3 2.5 82.5
+JSON $\times$ 28.3 25.3 5.0 82.3
+JSON $\checkmark$ 25.4 19.6 2.8 82.4
+[…]
+None $\checkmark$ 10.1 11.6 1.9 39.5
+Own $\checkmark$ 12.8 12.6 2.9 52.8
+Text $\checkmark$ 11.3 11.7 2.1 45.9
+JSON $\times$ 14.2 16.8 4.0 52.8
+JSON $\checkmark$ 15.5 14.5 15.3 53.3
+Table 1: Evaluation results using GPT3.5 (gpt35-16k).
+For each template, we experiment with different representations (Rep) and whether or not we include the style parameters (Style).
+We report the TAE F1 scores as calculated in §5, using ROUGE-L (R-L), METEOR (M), BLEU (B), and BERTScore (BERTS) as the similarity metrics.
 6.1 Results of automatic evaluation
-We report the results of our automatic evaluation in Table 2, computed the TAE scores as described in §5.
-As we can see from the results, first generating the SURe results in performance gains for most settings.
-We see that largest performance gains in the gpt35-16k experiments, indicating that our method has the largest benefits for simpler models.
-This is likely because more simple models have more difficulty parsing long unstructured documents, and the SURe provides a structured, filtered representation of the input document.
-The only setting we don’t see performance gains or similar performance is in the gpt4-32k setting on the poster template.
-Upon closer inspection, the references in this dataset are very verbose, averaging 391 tokens total.
-Our method produces generally less verbose posters, averaging 265 total tokens compared to 345 tokens produced by the baseline.
-We hypothesize that by editing the style parameters to include information about verbosity and length, we can improve performance on posters and leave exploration of this to future work.
-6.1.1 Importance of Structure
-Our method provides a template JSON to the LLM to generate the SURe, before passing the resulting JSON back to the LLM to generate the final document.
-A natural question is, how important is the JSON format?
-In other words, do we see performance gains as a result of the filtering of information, or is the structure provided by the JSON important as well?
-In order to test this question, we conduct an additional experiment in which we take the text within the SURe, discard the JSON structure, and pass the resulting body of text to the LLM to generate the final document.
-We randomly sample 100 documents each from LongSumm and Doc2PPT for the blog and poster generation tasks, respectively.
-We use the entirety of the Paper-Poster dataset, since it contains less than 100 examples.
-We use gpt35-16k for this experiment.
-The results of this experiment can be found in Table 3.
-As we can see from the results, passing the text of the SURe without the JSON structure decreases performance in most cases, with posters having the largest gap in performance, indicating that both the structure and content of the SURe are important for performance.
-Template
-SURe
-Format
-R-1 R-L
-Bert
-Score
-slides JSON 7.7 7.3 35.1
-Text 7.9 7.1 34.6
+In Table 1, we report the TAE F1 scores as described in §5, using ROUGE-L Lin (2004), METEOR Banerjee and Lavie (2005), BLEU Papineni et al.
+(2002), and BERTScore Zhang* et al.
+(2020) for the similarity measure.
+As seen in the results, by most measures, generating a JSON intermediate representation yields the best performance.
+We see that using the text representation generally degrades the performance over providing the structured JSON representation, indicating that structure is important for downstream performance in addition to abstractive filtering of information.
+Additionally, the text representation performs better than skipping the intermediate step altogether for both the poster and slide generation task, but not the blog generation task.
+This is likely because posters and slides have more inherent structure than blog posts, which can be relatively free-form.
+Finally, we see that allowing the model to choose its own representation format degrades performance over providing our pre-defined JSON structure.
+However, we see that in most cases, providing a representation generated without a JSON structure still performs better than skipping the intermediate generation step altogether (while maintaining the same style parameter setting).
+This indicates that even without a pre-defined structure, the intermediate step is still valuable for performance.
+Similarity Measure
+Model Rep. R-L M B BERTS
+Slides MS $\times$ 0.6 0.4 0.0 28.6
+$\checkmark$ 4.4 4.6 0.4 30.5
 […]
-Text 28.3 24.4 82.5
+$\checkmark$ 6.7 7.9 1.0 34.0
 […]
-Text 11.5 10.5 46.3
-Table 3: TAE scores comparing the use of the JSON representation of the SURe or passing the text generated in the SURe, using gpt35-16k.
-We see that in most cases, the JSON representation performs better, indicating the importantance of structure.
+$\checkmark$ 8.4 9.1 2.0 38.1
+[…]
+$\checkmark$ 21.7 16.2 1.7 81.3
+[…]
+$\checkmark$ 25.6 20.5 3.2 82.5
+[…]
+$\checkmark$ 25.8 20.2 3.1 82.7
+[…]
+$\checkmark$ 6.0 6.5 1.2 38.1
+[…]
+$\checkmark$ 10.4 11.0 1.5 50.7
+[…]
+$\checkmark$ 14.6 15.3 3.7 57.2
+Table 2: TAE F1 scores using Mistral-7b (MS), Mixtral (MX), and GPT4.
+We use ROUGE-L (R-L), METEOR (M), BLEU (B) and BERTScore (BERTS) as our similarity measures.
+For each template, we compare a JSON representation versus skipping the intermediate generation step (Rep), maintaining the same style parameters in both settings.
+Experiments with additional models.
+[…]
 Figure 3: Reasons annotators preferred each document.
-While annotators largely preferred documents generated with a SURe, the most common reasons for preference are better formatting and information content.
+While annotators largely preferred documents generated with an intermediate representation, the most common reasons for preference are better formatting and information content.
 We exclude the “Other” count as it was only selected once.
 6.2 Human evaluation
-We additionally conduct a human evaluation to test both our method and how well our evaluation framework aligns with human preference.
-We sample 100 documents each from Doc2PPT and LongSumm, and use the entirety of the Paper-Poster dataset.
-We present the annotators with 2 versions of each document, one generated with the SURe and one without.
-Both versions use gpt4-32k.
-The annotators are provided with the original paper and the intended document type (blog, slide deck, or poster), and are asked the following questions:
-  1. 1.
-    Which document do you prefer?
-  2. 2.
-    On a scale of 1-3, to what degree do you prefer your selection?
-  3. 3.
-    Why do you prefer your selection?
-For Question 3, the annotators are given a multi-select checklist of reasons why they might prefer one document over the other.
-The checklist included (1) the quality of the content, (2) the formatting, (3) the document style matching the intended document types, (4) the information represented in the document, and (5) other.
-The “other” option included a free text box.
-The full annotation instructions, including a description of the reasons provided and examples, can be found in §A.
-Each document was annotated by 3 different annotators.
-We employed 4 annotators from India, sourced via a third-party agency, to carry out the human evaluation of our task.
-A guideline document was provided to these evaluators, containing task-specific instructions, guidance, and several annotated examples.
-They were compensated at a rate of \$11.98 USD per hour for the total time that they spent working on the task, including a training round of annotation.
+After showing that LLMs benefit from intermediate structured representations in document transformations, we investigate whether our proposed evaluation framework aligns better with human judgment than previously proposed metrics.
+We sample 100 documents each from DOC2PPT and LongSumm, and use the entirety of the Paper-Poster dataset in this study.
+We present annotators with 2 versions of each document, one generated with the intermediate representation and one without.
+Both versions use gpt4-32k, as the best performing model.
+The annotators are provided with the original paper and the intended document type (blog, slide deck, or poster), and are asked the following questions: (1) Which document do you prefer?
+(2) On a scale of 1-3, to what degree do you prefer your selection?
+(3) Why do you prefer your selection?
+For question 3, annotators are also provided with a multi-select checklist of reasons for their preference: (1) quality of the content, (2) formatting, (3) document style matching the intended document types, (4) information represented in the document, and (5) other (along with a free text box).
+The full instructions, including the reasons provided and examples, can be found in Appendix §B.
+[…]
 Which method do humans prefer?
-The documents generated with a SURe were preferred by 82%, based on majority vote.
-71% of the documents generated with a SURe were unanimously preferred by the annotators.
-We calculate a annotator agreement score of 0.51 with Krippendorff’s alpha, indicating that while this is a subjective and specialized task, even non-expert annotators agree to a moderate degree.
+The documents generated with an intermediate representation were preferred by 82%, based on majority vote (71% unanimously).
+The annotator agreement score was 0.51 with Krippendorff’s alpha, indicating that while this is a subjective and specialized task, even non-expert annotators agree to a moderate degree.
 A visualization of the reasons the annotators preferred their selection can be found in Figure 3.
-We can see from the results that while annotators largely preferred the documents generated with a SURe, the most common reasons for preference are better formatting and better information content.
-This indicates that the structure provided by the SURe makes it easier for the model to format the final document well.
-Additionally, the SURe only includes the most salient information from the original text, resulting in higher quality of information content.
-Finally, we see a fairly even distribution across the different templatic views for the reasons of preference, indicating humans prefer the documents generated with the SURe across different document types.
+It can be seen that while annotators largely preferred the documents generated with an intermediate representation, the most common reasons for preference are better formatting and better information content.
+This indicates that the structure provided by the intermediate representation makes it easier for the model to format the final document well.
+Additionally, the intermediate representation only includes the most salient information from the original text, resulting in higher quality of information content.
+Finally, we see a fairly even distribution across different templatic views for the reasons of preference, indicating humans prefer the documents generated with the intermediate representation across different document types.
+Which metric correlates better with humans?
+We test whether our metric, as described in §5, correlates better with human preference compared to prior evaluation metrics in the literature.
+For each annotation, given the degree of the preference $d$ (Appendix §B Q.
+2) we convert value to a score $P(d)\rightarrow[1,2,3]$ if $d$ is slight, moderate, or strong, respectively.
+If the annotator prefers the document generated without an intermediate representation, we take $-P(d)$ instead.
+This allows us to measure if the metric captures directionality of preference along with degree.
+in parallel, we compute the automatic score $m$ for each metric, then calculate $S=m(\textrm{with rep})-m(\textrm{skip rep})$ where $m$ is the metric we are evaluating (e.g ROUGE).
+If a human annotator prefers a document generated without the intermediate step, we’d expect a good metric to assign a higher score to that document as well, resulting in both $S$ and $P(d)$ being negative (and positive in the opposite case).
+Using this intuition we assign an affinity score of a metric with respect to human evaluation as the Pearson correlation Freedman et al.
+(2007) of $S$ and $P(d)$.
+Since prior metrics are not designed to account for the structure of documents, we compute them by extracting only the text of both the generated and reference documents.
+The correlations with human judgement for each metric to its respective TAE variants can be found in Table 3.
+As we can see from the results, evaluations using our template adaptable framework correlate more highly with human judgement, except in the case of BERTScore.
+In the latter case the results are not statistically significant, and we hypothesize that the open-domain nature of BERT embeddings are poorly suited to represent the semantic similarity of scientific text.
 Metric PearsonR
-TAE ROUGE-1 21.4
-ROUGE-1 14.9
+ROUGE-L 14.5
+TAE ROUGE-L 19.7
 […]
-ROUGE-L 14.5
+TAE METEOR 25.2
 […]
-BERTScore 10.6*
+TAE BLEU 13.8
+[…]
+TAE BERTScore 5.4*
 Table 3: Correlation of evaluation metrics with human judgement.
 We compare each metric computed using the TAE framework versus the standard computation.
-*Indicates the correlation is not statistically significant ($p>0.01$)
-Which metric correlates better with human judgement?
-We use the human preference judgements to test if our metric, as described in §5, correlates more with human preference compared to prior evaluation metrics in the literature.
-For each annotation, given the degree of the preference $d$ we convert the model preference to a score $P$ as follows:
-$P(d)=\begin{cases}3,&d=\textrm{strong}\\ 2,&d=\textrm{moderate}\\ 1,&d=\textrm{slight}\end{cases}$
-If the annotator prefers the document generated without a SURe, we multiply $P(d)$ by $-1$. This allows us to measure if the metric captures both directionality of preference and degree.
-Finally, we compute the automatic score for each metric, then calculate $S=m(\textrm{with SURe})-m(\textrm{skip SURe})$, where $m$ is the metric we are evaluating.
-If the annotator prefers the document generated without the SURe, we’d expect a good metric to assign a higher score to that document as well, resulting in both $S$ and $P$ being negative.
-Conversely, we’d expect instances where the opposite were true to result in positive scores for both $S$ and $P$.
-Using these intuitions we assign an affinity score of a metric with respect to human evaluation as the Pearson correlation Freedman et al.
-(2007) of $S$ and $P$.
-Since prior metrics are not designed to account for the structure of the document, we compute them by extracting only the text of both the generated document and reference documents.
-The correlations with human judgement for each metric, computed standard compared to TAE variants can be found in Table 4.
-As we can see from the results, computing each evaluation metric using our template agnostic evaluation method correlates more highly with human judgement, except in the case of BERTScore.
-However, BERTScore is the only metric in which the correlation is not statistically significant, as it is likely simply a poor metric for this task.
-We hypothesize that BERTScore does not perform well on scientific data, since it is trained primarily on out of domain text.
+*Indicates the correlation is not statistically significant ($p>0.01$).
 7 Conclusion
-In many domains, people choose to disseminate information across different modalities and formats, allowing for better communication of information to broader audiences.
-Past work has focused on developing methods to generate different templatic views of documents, tailoring the method to each template.
-In this paper, we introduced an LLM-powered method to unify the generation of templatic views of documents by first generating a Structured Unified Representation (SURe) of the input document, representing the underlying knowledge of the templatic views.
-We showed that our method results in improvements in performance, especially for a simpler model.
-We additionally introduced Template Agnostic Evaluation (TAE), a flexible precision-recall framework for automatic evaluation that is template agnostic, allowing for easy integration of existing evaluation metrics into a unified system.
-The framework is capable of jointly accounting for and measuring multiple facets of the output document including the quality of its content, its underlying structure and ordering, as well as its length.
-Finally, we conducted a human evaluation and showed that annotators prefer the documents generated using the SURe 82% of the time.
-We additionally showed that our framework of evaluation correlates more highly with human preference than standard evaluation metrics.
-The use of templatic views of documents exists in many domains (e.g.
-creating a cover letter and personal website based on a resume), and in future work we will consider other domains of documents, as well as more complex scenarios, such as many-to-many document generation.
-Our work is a step towards a larger goal of generalized document generation.
-Limitations
-Although our methods are not domain specific, we only evaluated our methods in the scientific domain, due to the availability of public data.
-Further research is required to evaluate the performance of our methods in other domains.
-Additionally, our method is limited to textual content.
-Future work should explore incorporating multi-modal models and content generation.
-Finally, we evaluated our methods on closed source models, which may affect the reproducibility of our results.
-Appendix A Annotation Instructions
-A.1 Questions
+In many domains, people choose to disseminate information across different modalities and formats for better communication to broader audiences.
+We proposed a unified view of document transformation and evaluation.
+We showed that LLMs are capable of templatic document generation with minimal supervision, and that a structured, intermediate representation can improve performance, particularly for smaller models.
+We also introduced a flexible precision-recall framework for automatic evaluation that easily integrates existing evaluation metrics into a unified system and allows for comparison across diverse datasets without additional task specific metric design.
+Finally, we conducted a human evaluation and showed that annotators prefer the documents generated using the intermediate representation 82% of the time and that our evaluation framework correlates better with human preference than standard evaluation metrics.
+8 Limitations
+Although our methods are not domain specific, we only evaluated them in the scientific domain, due to the availability of public data.
+Additionally, our framework is limited to textual content.
+In future work we plan to explore the application of our unified framework for generation and evaluation on document views in other domains, as well as incorporating multi-modal models and content generation.
+Finally, it is possible that some of our test data has leaked into the training data of the models with which we experimented.
+This limitation is not unique to our work and exists for our baselines in addition to our methods.
+9 Ethics
+The potential risks of our work are similar to those of other work in downstream applications of LLMs.
+LLM generated documents can potential generate copy-righted material Carlini et al.
+(2020), personally-identifiable information Lukas et al.
+(2023), or factually incorrect text Manakul et al.
+(2023).
+The use of LLMs to generate documents may violate some academic dishonesty policies Zdravkova et al.
+(2023).
+Our system is intended to be used in collaboration with human writers.
+Users should edit the generations, checking for factual inconsistencies and other potential errors.
+Our work is intended to save users time that might be spent repeating information across multiple documents, so they can focus on content creation.
+Therefore, we believe the benefits of our work outweigh the potential risks.
+Appendix A Prompt details
+The prompts and intermediate representation template used can be found in Table 4 and Figure 4, respectively.
+We note that the specific structure provided to the prompt is not inherent to our method, and a different structure could be provided depending on the input document and domain.
+For the tasks we evaluate in this paper, we use the following style parameters:
+  • Slides: “Slides should include a title page. Following slides should contain an informative slide title and short, concise bullet points. Longer slides should be broken up into multiple slides.”
+  • Posters: “Posters should include a title section at the top. Each panel should include a heading and short, concise bullet points of the most important take-aways from that section.”
+  • Blogs: “Blogs should include paragraphs introducing the topic, a summary of the input document, and important takeaways.
+The blog should be more readable to a general audience than the input document.”
+Figure 4: Template of the intermediate representation provided to the prompts in Table 4.
+Appendix B Annotation Instructions
+B.1 Questions
 Question 1 – Which document do you prefer? In this question, you are asked to choose which document version you prefer. Some examples of qualities you may use to decide your preference include:
@@
 Not all criteria must be met, and you may use other relevant criteria to make your decision.
-You are not rating the document for factual correctness, and only need to refer to the corresponding scientific article if it will aid in making your preference.
+You are not rating the document for factual correctness, [^6] The annotators are non-experts and do not have the background to determine factual correctness of scientific information.
+Instead, they are encouraged to use the original paper to understand if the information presented in the documents represent the information in the paper, to the best of their understanding.
+and only need to refer to the corresponding scientific article if it will aid in making your preference.
 You can answer this question with either Document A or Document B.
@@
     Other (free text)
-A.2 Edge cases
+B.2 Edge cases
 For most edge cases, it is up to your discretion on how to best handle the case. However, below are a few examples of how you could consider certain edge cases:
@@
 Since neither document matches the style or formatting of the intended document type, you could consider other criteria, such as quality of content or the information represented.
-Template Temp R-1 R-L BertScore
-slides 0 7.7 7.3 35.1
-0.25 7.8 7.2 35.4
-0.5 7.6 7.0 36.4
-0.75 7.5 7.0 35.5
-1 8.0 7.4 35.8
+Simularity Measure
+Temp R-L M B BERTS
+Slides 0.0 7.3 8.2 1.6 35.1
+0.25 7.2 8.3 1.5 35.4
+0.5 7.0 8.3 1.5 36.4
+0.75 7.0 8.0 1.2 35.5
+1.0 7.4 8.2 1.4 35.8
 […]
-0.25 27.2 25.5 82.6
-0.5 28.0 26.2 82.8
-0.75 27.2 25.4 82.6
-1 26.6 24.8 82.7
+0.25 25.5 19.8 2.7 82.6
+0.5 26.2 20.9 3.2 82.8
+0.75 25.4 19.9 2.7 82.6
+1.0 24.8 19.9 2.6 82.7
 […]
-0.25 14.2 13.0 53.1
-0.5 13.6 12.5 52.2
-0.75 13.2 12.0 50.8
-1 12.6 11.6 50.3
-Table 5: Results of the temperature hyperparameter experiments. There seems to be little consistency in performance as related to temperature across different templates.
-Appendix B Temperature Experiments
+0.25 13.0 14.8 3.4 53.1
+0.5 12.5 14.0 2.7 52.2
+0.75 12.0 13.9 3.0 50.8
+1.0 11.6 11.9 2.4 50.3
+Table 5: Results of the temperature hyperparameter experiments. We use ROUGE-L (R-L), METEOR (M), BLEU (B) and BERTScore (BERTS) as our similarity measures.
+Appendix C Temperature Experiments
 We experiment with the temperature of the generations to see how temperature affects performance.
-Similar to the textual representation experiments, described in §6.1.1, we randomly sample 100 documents each from LongSumm and Doc2PPT for the blog and poster generation tasks, respectively.
+We randomly sample 100 documents each from LongSumm and Doc2PPT for the blog and poster generation tasks, respectively.
 We use the entirety of the Paper-Poster dataset, since it contains less than 100 examples.
@@
 As we can see from the results, there seems to be little consistency across the different types in which temperature performs the best.
+Appendix D Longsumm Blind Test Set Results
+We submit the final documents from GPT4, the best performing model overall, to the Longsumm blind test set evaluation.
+We compare the documents generated with and without the intermediate step.
+We see that without the intermediate representation we get a Rouge-1 score of 46.8 while the results generated without the intermediate representation received a Rouge-1 score of 46.4.
+We note that this blind test set of 22 papers is significantly smaller than the evaluation data (505 papers) we used in the main body of this paper.
+Despite not designing a task specific method, we place second on the leaderboard, showing the powerful capabilities of LLMs in long document generation.
+Appendix E Example Outputs
+We provide examples of the outputs generated with and without the intermediate representation below.
+The documents in all examples are generated with GPT4 (gpt4-32k).
+Figure 5 includes example slide generations, Figure 6 includes example blog generations, and Figure 7 includes example poster generations.
+(a) Document generated without intermediate representation. This example is not cropped.
+(b) Document generated with intermediate representation. This example is cropped for space and includes an additional 4 slides that are not included for space.
+Figure 5: The above documents are example slides generated by GPT4 (gpt4-32k) with and without the intermediate representation.
+We can see that without the intermediate step, the model did not generate a true slide deck.
+(a) Document generated without the intermediate representation. This example is not cropped.
+(b) Document generated with the intermediate representation. This example is cropped for space and includes an additional page of text.
+Figure 6: The above documents are example blog posts generated by GPT4 (gpt4-32k) with and without the intermediate representation.
+We can see that without the intermediate representation, the model did not properly format the LaTeX file for compilation.
+(a) Document generated without the intermediate representation. This example is cropped for space and includes an additional 3 slides.
+(b) Document generated with the intermediate representation. This example is cropped for space and includes an additional 4 slides.
+Figure 7: The above documents are example posters generated by GPT4 (gpt4-32k) with and without the intermediate representation.
+We found that GPT4 often generates slide decks in place of posters.
+We can see that the document generated without the intermediate representation contains more verbose panels and includes less formatting.